discuss the continuity/uniform continuity/Lipchitz continuity and differentiability of the functions |x| on [βˆ’1,1].


1
Expert's answer
2021-09-27T14:31:49-0400

ANSWER.Function "f(x)=|x|" is Lipschitz continuous, therefore, uniformly continuous, hence continuous on [-1,1]. The function is not differentiable at all points of the [-1,1] (At the point x=0 the function is not differentiable).

EXPLANATION.

Function "f(x)=|x|" is Lipschitz continuous, because for all "{ x }_{ 1 },{ x }_{ 2 }\in \left[ -1,1 \right] :\ \left| f({ x }_{ 1\ })-f\ ({ x }_{ 2 }) \right| =\left| \left| { x }_{ 1\ } \right| -\left| { x }_{ 2\ } \right| \right| \ \le \left| { x }_{ 1 }-{ x }_{ 2 } \right|" (1).

Since for any "\u03b5>0" and "\u03b4=\u03b5" for any "{ x }_{ 1 },{ x }_{ 2 }\in \left[ -1,1 \right]" such that "\left| { x }_{ 1 }-{ x }_{ 2 } \right| <\delta \quad"from (1) it follows "\left| f({ x }_{ 1\ })-f\ ({ x }_{ 2 }) \right| =\left| \left| { x }_{ 1\ } \right| -\left| { x }_{ 2\ } \right| \right| < ε" , then "f" is uniformly continuous.

"f" is differentiable on "[-1,0)" and "(0,1]" , because

"f(x)=\begin{cases} -x,-1\le x<0 \\ 0,\quad x=0 \\ x,\quad 0<x\le 1 \end{cases}, f'(x)=\begin{cases} -1,-1\le x<0 \\ 1\ ,\quad 0<x\le 1 \end{cases}" .

Since "{ f }_{ - }^{ ' }(0)=\lim _{ x\rightarrow { 0 }^{ - } }{ \frac { f(x)-f(0) }{ x } } =\lim _{ x\rightarrow { 0 }^{ - } }{ \frac { -x-0 }{ x } =-1,\quad { f }_{ + }^{ ' }(0)= } \lim _{ x\rightarrow { 0 }^{ + } }{ \frac { f(x)-f(0) }{ x } } =\lim _{ x\rightarrow { 0 }^{ + } }{ \frac { \ x-0 }{ x } =\ 1 }"

then the left derivative is not equal right derivative. So , the function is not differentiable at the point "x=0."